OpenStudy (anonymous):

Solve: 8x − 10 + 18 = 24 4 −2 6.5 2

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

8x − 10 + 18 = 24 8x+8=24 8x=24-8 8x=16 x=16/8 x=2